OpenStudy (anonymous):

Population t years from now: p(t)=4t+6t^(3/2)+200 a) At what rate will the population be changing 4 years from now? b) what is the estimated population 4 years from now?

OpenStudy (anonymous):

dp(t) /dt = 4 + (6*3t^0.5)/2 dp(t) /dt = 4 + 9t^0.5 a) plug t = 4 4+ 9*2 = 22 b) p(4) = 4*4 + 6*4^(3/2) + 200 = 216 + 6 * 8 = 264